OpenStudy (anonymous):

Which line is a graph of the equation: –2x + 5y = 10? VHS_ALG_S1_07_L203_L303_LQ4_graphi.gif A. line a B. line b C. line c D. line d

OpenStudy (solomonzelman):

–2x + 5y = 10 translate into slope intercept form. add 2x to both sides 5y=2x+10 and divide both sides by 5 y=2/5x-10 plug in values for x, solve for y. You should find at least 2 (x,y) coordinates. So plug any number for x, at least twice. Connect the points with the line. Once you get the graph, you will know.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

there it is srry

OpenStudy (solomonzelman):

Do what I told you. You will find out.

OpenStudy (anonymous):

what line is it